# Intensity-Efficient Allocations

### v5

@@ 2 Notation, Preliminaries and Measurement-Theoretic Foundations
A relation $\dot{\succsim}_{l}$ that is represented by $s_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widehat{s}_{l}$ of $s_{l}$ satisfies
$\widehat{s}_{l}(a,c)$ $=$ $\widehat{s}_{l}(a,b)+\widehat{s}_{l}(b,c)$ (7)
Proof.
Upon viewing $\widehat{s}_{l}$ as a general function of two variables, (7) is known as Sincov’s functional equation, whose solution is some $f:A\rightarrow\mathbb{R}$, unique up to an additive constant, such that $\widehat{s}_{l}(a,b)\equiv f(a)-f(b)$ (Aczél, 1966). This establishes the “if” part. The “only if” part is immediate upon defining $\widehat{s}_{l}(a,b):=u_{l}(a)-u_{l}(b)$ for the postulated function $u_{l}$ with the requisite representation property. ∎
The normalization of $s_{l}$ introduced next is important for the rest of this paper.
@@
A preference intensity function $s_{l}$ represents $\dot{\succsim}_{l}$ canonically if
$s_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
@@
The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
@@
Proof.
Let $s_{l}$ be an arbitrary preference intensity function for $\dot{\succsim}_{l}\in\mathcal{I}$. Define the level set of $s_{l}$ at $(a,b)\in A\times A$ by $\mathbf{s}_{l}[a,b]:=\{(a^{\prime},b^{\prime})\in A\times A:s_{l}(a,b)=s_{l}(a^{\prime},b^{\prime})\}$. Further, define $r_{l}:A\times A\rightarrow\mathbb{R}$ by
By construction, $r_{l}$ is canonical and $r_{l}(a,b)\geq r_{l}(c,d)\Leftrightarrow s_{l}(a,b)\geq s_{l}(c,d)$. ∎
@@
This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$. It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies.
As we elaborate in the next section, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$. In particular, the only non-trivial intensity-equivalence class that exists when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
@@
$n=$ orders intensity relations intensity relations profiles
@@
Finally, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit, respectively, a (pseudo-cardinal) utility-difference and an ordinal representation when $n\leq 5$, and was produced with Minizinc (Nethercote et al., 2007; Gecode Team, 2006).
3 The Comparability Postulate
@@
Interpersonal Comparisons of Ordinal Intensities
@@
Towards defending this new concept we first recall that preference intensities at the level of the individual decision maker are not assumed here to be quantifiable beyond the level of an ordinal ranking. Yet some information about the generally different welfare effects of having $a$ or $b$ vs $c$ or $d$ is available here. Suppose, in particular, that agent $l$ prefers $a$ to $b$ to $c$ to $d$, and the first over the second more than the third over the fourth. This is equivalent to saying that her intensity difference at pair $(a,b)$ is greater than at $(c,d)$. Consider agent $m$ next, and suppose for simplicity that the same applies to him too: $s_{m}(a,b)>s_{m}(c,d)$. Considering this, and that both have preferences and intensities over the same set of alternatives, how might (10) be interpreted?
@@
It is clearly a statement of fact that the intensity difference within pair $(a,b)$ is ranked higher by agent $l$ in her intensity ordering than the corresponding difference within pair $(c,d)$ is ranked by agent $m$ in his. Now recall that, by the Strictness assumption, $l$ and $m$ have the same number of possible rank positions in their respective intensity orderings which pairs $(a,b)$ and $(c,d)$ can occupy. In other words, the agents’ ordinal intensity scales coincide: if $(a,b)$ and $(c,d)$ are ranked 5th and 8th in $l$’s and $m$’s scales, respectively, then they are ranked 5th and 8th out of the same $k$ possible positions. As we mentioned in the Introduction, this logic resembles the way scoring rules such as the well-known Borda-count (Borda, 1781) normalization of ordinal utilities or scoring-like (see d’Aspremont, 2018) rules such as the relative-utilitarian (Dhillon and Mertens, 1999) normalization of cardinal utilities perform interpersonal comparisons in their respective environments.
@@
The question now emerges: Should the different agents’ intensity orderings be treated equally by the social planner? While equal treatment may or may not be an appropriate approach to follow in practice, depending on what else is known about agents and the problem at hand, since the intensity orderings are assumed here to encode all available welfare-relevant information, treating them in any way other than equal would call for a justification that appears elusive. Against this background, we view the suggested interpretation of (10) as a reasonable starting point for interpersonal comparisons in this informational environment.
@@
On the analytical side, another relevant question also arises: in the special case where the ordinal intensity ordering of every agent is defined by how the pairwise differences of a cardinal utility function are ordered, and these differences in turn define a strict intensity profile, does interpersonal comparisons of ordinal intensities reduce to the interpersonal comparability of utility differences that is required by utilitarian aggregation? As the next example shows, the answer is no.
@@
Let $n=3$ and consider the utility profile $U:=\big{(}(53,38,9),(46,33,21),$ $(20,30,50)\bigr{)}$ over $A=\{a_{1},a_{2},a_{3}\}$, where each $u_{l}$ in $U$, $l\leq 3$, is constructed so its values add up to 100 (this is without loss). It is readily seen that the intensity profile defined by the intra-personal utility differences in $U$ is in $\mathcal{\widehat{S}}$. Also, defining $\dot{\succ}_{i}$ by $(a_{i},a_{j})\dot{\succ}_{l}(a_{k},a_{m})$ $\Leftrightarrow$ $u_{l}(a_{i})-u_{l}(a_{j})>u_{l}(a_{k})-u_{l}(a_{m})$, one observes that $(a_{2},a_{3})\,\dot{\succ}_{1}\,(a_{1},a_{2})$ and $(a_{1},a_{2})\,\dot{\succ}_{2}\,(a_{2},a_{3})$ hold, hence $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$ too. However, $u_{1}(a_{1})-u_{1}(a_{2})>u_{2}(a_{1})-u_{2}(a_{2})$ is also true. $\blacklozenge$
@@
4 Intensity Dominance and Intensity-Efficient Allocations
4.1 Definitions
Since the number of agents and items are assumed to coincide, an allocation of the $n$ goods is a permutation on $A$. The set of all allocations is denoted by $\mathcal{A}$. As we discussed in the Introduction, we are interested in the assignment of the $n$ objects in $A$ to the $n$ agents in $V$ in a way that satisfies Pareto-efficiency but also improves upon it in delivering a narrower set of acceptable allocations that satisfy some additional desideratum that is appealing from a distributive-justice perspective and is operational in the present analytical environment where information on ordinal preference intensities and interpersonal comparisons thereof is assumed to be available. To this end, we proceed with the introduction of the following novel notions of dominance and efficiency.
@@ 4.1 Definitions
Let $S=\big{(}\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n}\bigr{)}\in\widehat{\mathcal{S}}$ be a strict intensity profile and $s$ its canonical representation. Given allocations $x$ and $y$, the former is said to intensity-dominate the latter if, for every pair of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$ it holds that $s_{i}(x_{i},x_{j})\geq s_{j}(y_{j},y_{i})$, and there is at least one pair $(i,j)$ where this inequality is strict.
@@
An allocation is intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ if it is Pareto-efficient at the induced preference profile $P_{S}\in\mathcal{P}^{V}$ and is not intensity-dominated by another Pareto-efficient allocation at $P_{S}$. It is intensity-dominant if it intensity-dominates every other Pareto-efficient allocation at $P_{S}$.
@@ 4.2 Intensity-Efficient and Utilitarian Allocations Are Distinct
To enable a comparison between the concept introduced above and the benchmark Pareto-efficiency refinement afforded by classical utilitarianism, we assume throughout this subsection that, in addition to satisfying the strictness condition $(\mathbf{5_{\mathit{D}}})$, each $\dot{\succsim}_{l}$ is utility-difference representable as in (3) under some cardinal utility index $u_{l}$. This is the case, for example, when $u_{l}:A\rightarrow\mathbb{R}$ is the utility function associated with an expected-utility representation of preferences over lotteries over $A$ and the intensity order defined by the $u_{l}$-differences is strictly convex or strictly concave relative to the agent’s induced (strict) preferences. With this assumption in place, we finally invoke the Cardinal Unit Comparability (Roemer, 1996; Baccelli, 2023) informational basis to social welfare. More specifically, we assume that: (i) $u_{l}:A\rightarrow\mathbb{R}$ is unique up to a positive affine transformation and features utility differences that define an intensity relation $\dot{\succsim}_{l}$ that abides with $(\mathbf{5_{\mathit{D}}})$; (ii) interpersonal utility-difference comparisons are possible under such transformations of any utility profile $U=(u_{1},\ldots,u_{n})$ where each $u_{l}$ is of this kind, as long as the scale/multiplicative parameter of such transformations is the same for all $u_{l}$. As is well-known, this informational basis is sufficient [^1] The reason why it is not necessary is closely related to the complex uniqueness properties of general utility-difference representations in the sense of (3). See Bossert (1991) for more on this. for the classical utilitarian social welfare functional—recalled below—to be meaningful.
@@
Proof.
Consider the strict utility profile of Example 1. Observe that the Pareto-efficient allocations are $(a_{1},a_{2},a_{3})$ and $(a_{2},a_{1},a_{3})$, and that the former allocation is the utilitarian one. Yet this is intensity-dominated by the latter since $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$. ∎
This potential divergence between intensity efficiency and the hallmark cardinal-welfarist allocation criterion is important because it shows that the hereby proposed notion is indeed distinct even under the richest possible informational assumptions. Furthermore, while perhaps not obvious ex ante, as is clarified by Example 1 and the discussion at the end of in Section 3, the reason why such a divergence may occur is that the classical utilitarian criterion is guided by the utility differences across agents. By contrast, in this special environment where intensity orderings are defined by the relative ordering rather than the absolute magnitude of utility differences, the Interpersonal Comparisons of Ordinal Intensities postulate upon which the intensity efficiency criterion is built compares the relative ranking of those differences across agents’ intensity orderings.
@@
Proof.
Under the postulated conditions, suppose to the contrary that $x(U)$ is a utilitarian but not intensity-efficient allocation. Then, given that agent $l$’s intensity relation $\dot{\succsim}_{l}$ is defined by $(a,b)\,\dot{\succsim}_{l}\,(c,d)\Leftrightarrow u_{l}(a)-u_{l}(b)\geq u_{l}(c)-u_{l}(d)$, from the definition of intensity-dominance and intensity-efficiency there exist at least one pair of agents $(i,j)$ and of objects $(a,b)$ such that $x_{i}(U)=a$, $x_{j}(U)=b$ and $s_{j}(a,b)>s_{i}(a,b)$. By balancedness, this implies
$u_{j}(a)-u_{j}(b)$ $>$ $u_{i}(a)-u_{i}(b)$ (13)
Consider allocation $x^{\prime}$ that is identical to $x(U)$ except that $x^{\prime}_{i}=b$ and $x^{\prime}_{j}=a$. From (13) we get $u_{j}(a)-u_{i}(a)>u_{j}(b)-u_{i}(b)$. This shows that the utilitarian benefit from the $i\leftrightarrow j$ swap in the transition from $x(U)$ to $x^{\prime}$, $u_{j}(a)-u_{i}(a)$, exceeds the corresponding loss, $u_{j}(b)-u_{i}(b)$. It follows that $\sum_{l=1}^{n}u_{l}(x^{\prime}_{l})>\sum_{l=1}^{n}u_{l}\big{(}x_{l}(U)\bigr{)}$, which contradicts the postulated utilitarian-efficiency of $x(U)$. ∎

### v6

@@ 2 Measurement-Theoretic Foundations
A relation $\dot{\succsim}_{l}$ that is represented by $\widehat{s}_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widetilde{s}_{l}$ of $\widehat{s}_{l}$ satisfies
$\widetilde{s}_{l}(a,c)$ $=$ $\widetilde{s}_{l}(a,b)+\widetilde{s}_{l}(b,c)$ (7)
The proof of this and most other results appear in the Appendix.
The normalization of $\widehat{s}_{l}$ introduced next is important for the rest of this paper.
@@
A preference intensity function $\widehat{s}_{l}$ represents $\dot{\succsim}_{l}$ canonically if
$\widehat{s}_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
@@
We will denote by $s_{l}$ the canonical intensity function that represents $\dot{\succsim}_{l}$. The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
@@
This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$. It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies. As we elaborate in Section 3, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$. In particular, the only non-trivial intensity-equivalence class when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$. For illustration, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit,respectively, a (pseudo-cardinal) utility-difference representation and an ordinal representation when $n\leq 5$.
@@
$n$ orders intensity relations intensity relations profiles
@@
Source: author’s computations with the Minizinc constraint-programming tool (Nethercote et al., 2007).
We conclude this section by commenting on possible sources of intensity information. Section 2 in Gerasimou (2021) discusses various non-choice variables that, according to different literatures, might be thought of as having intensity-revealing potential. Prominent examples include willingness to pay; 50/50 lottery choices; survey-based ratings; and response times. Section 5 of that paper, moreover, shows how such data can be analysed alongside choices by means of otherwise standard revealed-preference theoretic techniques, providing testable foundations for the hypothesis that an agent’s overall behaviour complies with utility maximization and ordinal intensity revelation. The present paper, in contrast, operates within the environment of the allocation problem, where agents are traditionally assumed to submit their full ordinal or cardinal preference reports directly. This is different from a lab, field or other environment that might, in principle, provide the analyst with agents’ choices at several menus at different points in time, together with intensity-revealing information at those menus. As we elaborate formally in Section 6, for this specific environment we assume that agents’ ordinal preferences are available, and that their ordinal intensities are elicited by asking them directly—and conditional on their submitted preferences—the kinds of simple binary questions that were mentioned earlier.
3 Interpersonal (Non-)Comparability of Ordinal Intensities and Intensity-Rank Scores
@@
Assumption 1 (for Sections 4 & 5)
@@
Towards defending this assumption—whose importance, we stress, is interpretational—recall first that an individual’s preference intensities are not assumed to be quantifiable beyond the level of an ordinal ranking. Yet some information about the generally different welfare effects of having $a$ or $b$ vs $c$ or $d$ is available. Suppose, in particular, that $a\succ_{l}b\succ_{l}c\succ_{l}d$, and that the agent prefers $a$ to $b$ more than she does $c$ to $d$. This is equivalent to saying that her intensity difference at pair $(a,b)$ is greater than at $(c,d)$. Consider agent $m$ next, and suppose for simplicity that the same applies to him too: $s_{m}(a,b)>s_{m}(c,d)$. Given this, and given also that both agents have preferences and intensities over the same set of alternatives, how might (9) be interpreted?
@@
We know that the intensity difference within pair $(a,b)$ is ranked higher by agent $l$ in her intensity ordering than the corresponding difference within pair $(c,d)$ is ranked by agent $m$ in his. By the Strictness assumption, moreover, $l$ and $m$ have the same number of possible rank positions in their respective intensity orderings, which pairs $(a,b)$ and $(c,d)$ could possibly occupy. In other words, the agents’ ordinal intensity scales coincide: if $(a,b)$ and $(c,d)$ are ranked 5th and 8th in $l$’s and $m$’s scales, respectively, then they are ranked 5th and 8th out of the same $k$ possible positions. As we mentioned in the Introduction, this logic resembles the way scoring rules such as the well-known Borda-count (Borda, 1781) normalization of ordinal utilities or scoring-like (see d’Aspremont, 2018) rules such as the relative-utilitarian (Dhillon and Mertens, 1999) normalization of cardinal utilities perform interpersonal comparisons in their respective environments. Importantly, neither of these social welfare functions formally requires interpersonal utility comparisons in order for the respective social outcome to be immune to permissible transformations of the agents’ utility functions. Clearly, this is so because both welfare criteria are defined in terms of specific normalizations which reduce, respectively, any ordinally or cardinally equivalent utility function that represents an agent’s preferences into the same normalized ordinal or cardinal utility function.
At the same time, however, applications of the Borda and relative-utilitarian normalizations do introduce interpersonal utility comparisons. We elaborate on this in the case of the Borda method, which is conceptually closer to our setting. This takes agent $l$’s preference relation $\succ_{l}$ and defines the Borda score of every item $a$ in $A$ by
$b_{l}(a)$ $:=$ $|\{a^{\prime}\in A:a\succ_{l}a^{\prime}\}|,$ (10)
i.e. by the number of alternatives that $l$ considers inferior to $a$. Once these normalised utility functions are constructed for all agents, the Borda social welfare function maps any such profile to a social ordering that prioritises items with higher aggregate Borda scores. Although interpersonal comparability assumptions are not part of the axioms that characterize the Borda method, the motivation for its very existence does implicitly invoke such comparisons. This is articulated clearly in Maskin (2025, f.17), as follows: “Notice that the Borda count implies a particular way of making interpersonal comparisons, e.g., if individual 1 ranks $x$ two positions above $y$, that preference is exactly canceled by two individuals who rank $y$ one position above $x$. Observe, however, that none of our axioms speaks directly to interpersonal comparisons at all; such comparisons are an emergent property of the joint imposition of the axioms”.
@@
Returning to Assumption 1, we note the analogy between the canonically-normalized values that possibly different pairs of items may occupy in the intensity orderings of different agents, and the possibly distinct positions (Borda scores) of items across agents’s preference orderings. However, unlike the additional cancellation assumptions that are implicit in the aggregation of Borda scores, Assumption 1 merely requires that an ordinal comparison be made about $l$’s intensity between $a$ and $b$ vs $m$’s intensity between $c$ and $d$. The question now emerges: Assuming that it is desirable in the first place for this additional information that is contained in the different agents’ ordinal intensity orderings to be accounted for, should these be treated equally by the social planner? While equal treatment may or may not be an appropriate approach to follow in practice, depending on what else is known about agents and the problem at hand, since the intensity orderings are assumed here to encode all available welfare-relevant information, accepting their relevance but treating them in any way other than equal would call for a justification that appears elusive. Against this background, we view the suggested interpretation of (9) as a reasonable starting point for interpersonal comparisons of ordinal intensities in this informational environment.
@@
Another relevant question arises, now on the analytical side: in the special case where the ordinal intensity ordering of every agent is defined by how the pairwise differences of a cardinal utility function are ordered, and these differences in turn define a strict intensity profile, does interpersonal comparability of ordinal intensities in the sense of Assumption 1 reduce to the interpersonal comparability of utility differences that is required by utilitarian aggregation? As the next example shows, the answer is no.
@@
Let $n=3$ and consider the utility profile $U:=\big((53,38,9),(46,33,21),$ $(20,30,50)\bigr)$ over $A=\{a_{1},a_{2},a_{3}\}$, where each $u_{l}$, $l\leq 3$, is constructed so its values add up to 100 (this is without loss). It is readily seen that the intensity profile defined by the intra-personal utility differences in $U$ is in $\mathcal{\widehat{S}}$. Also, defining $\dot{\succ}_{i}$ by $(a_{i},a_{j})\dot{\succ}_{l}(a_{k},a_{m})$ $\Leftrightarrow$ $u_{l}(a_{i})-u_{l}(a_{j})>u_{l}(a_{k})-u_{l}(a_{m})$, we observe that $(a_{2},a_{3})\,\dot{\succ}_{1}\,(a_{1},a_{2})$ and $(a_{1},a_{2})\,\dot{\succ}_{2}\,(a_{2},a_{3})$, hence $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$ too. Yet $u_{1}(a_{1})-u_{1}(a_{2})>u_{2}(a_{1})-u_{2}(a_{2})$ is also true. $\blacklozenge$
@@
Although the analysis of Section 4 makes use of Assumption 1222In fact, it uses a weaker version where interpersonal comparisons are made when $(a,b)=(c,d)$ in (9)., Section 5 revolves around concepts that become meaningful under a richer interpersonal comparability assumption on agents’ ordinal intensities. Before stating it we first use an agent’s canonical intensity function to introduce the intensity-rank score (or intensity rank) of item $a$ by
$r_{l}(a)$ $:=$ $\max_{a^{\prime}\in A}s_{l}(a,a^{\prime})$ (11)
In Lemma 5 (see Appendix) we show that $r_{l}$ is an ordinal utility function for $l$’s preferences that are induced by $\dot{\succsim}_{l}$. Recall now that, under Assumption 1, the comparison $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ is interpreted as suggesting that the intensity difference between $a$ and $b$ is at least as large for agent $l$ as the respective difference between $c$ and $d$. Logically, then, $r_{l}^{S}(a)>r_{l}^{S}(b)$ means that the maximum intensity difference between $a$ and some alternative that is inferior to it exceeds the maximum such difference that corresponds to $b$. In other words, the intensity-rank score of $a$ reflects the highest position in $l$’s intensity ordering where $a$ finds itself being preferred to another item. By its definition, therefore, and unlike the Borda score $b_{l}(a)$, this new score contains information both about how many alternatives are worse than $a$ in $l$’s preference ranking and about the intensity difference between $a$ and its inferior alternatives.
Assumption 2 (for Section 5)
$r_{l}(a)$ $>$ $r_{m}(b)$ (12)
is interpreted as suggesting that agent $l$ prefers $a$ more than agent $m$ prefers $b$.
To motivate this assumption and connect it to the previous one, suppose $r_{l}(a)\equiv s_{l}(a,a_{w})$ and $r_{m}(b)\equiv r_{m}(b,b_{w})$, where $a_{w}$ and $b_{w}$ are the bottom-ranked items in $A$ for agents $l$ and $m$, respectively. Let $r_{l}(a)>r_{m}(b)$. By definition, this is equivalent to $s_{l}(a,a_{w})>s_{m}(b,b_{w})$. Under Assumption 1, this means that $l$ prefers $a$ to their worst possible item more than $m$ prefers $b$ to their own worst possible item. Assumption 2, then, makes this last statement the definition of an interpersonal utility comparison between $a$ and $b$ for agents $l$ and $m$. Utility functions defined by Borda scores are onto the same consecutive integral range $\{0,1,\ldots,n-1\}$. By contrast, the extended—but still common across agents—range of utility functions defined by intensity-rank scores is $\{0,1,\ldots,{{n}\choose{2}}\}$. Those functions are obviously not onto this range, and their images are generally distinct, although, by definition, they all have the same minimum and maximum values. We will elaborate further on the relation between $b_{l}$ and $r_{l}$ in Section 5.
We conclude this section by stressing that every welfare-theoretic concept that we introduce in the sequel relies on the position where a pair of alternatives lies within an agent’s intensity ranking. Thus, by introducing canonically normalized intensity functions at this early stage we assign a specific name and piece of notation for these score-like positions. Crucially, this means that the welfare concepts in later sections are invariant with respect to arbitrary monotonic transformations of any intensity functions that one may choose to represent the agents’ ordinal intensities by, because—similar to the Borda and relative utilitarian functions—these are automatically translated into canonically normalized values of ordinal intensity functions that are onto the same range for all agents.
4 Intensity-Dominance and Efficiency
4.1 Intensity-Dominance and Efficiency
Since the number of agents and items are assumed to coincide, an allocation of the $n$ goods is a permutation on $A$. The set of all allocations is denoted by $\mathcal{A}$. As we discussed in the Introduction, we are interested in the assignment of the $n$ objects in $A$ to the $n$ agents in $N$ in a way that satisfies Pareto-efficiency but also improves upon it in delivering a narrower set of acceptable allocations that satisfy some additional desideratum that is appealing from a distributive-justice perspective and is operational in the present analytical environment where information on ordinal preference intensities and interpersonal comparisons thereof is assumed to be available. To this end, we proceed with the introduction of the following novel notions of dominance and efficiency.
@@ 4.1 Intensity-Dominance and Efficiency
Let $S=\big(\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n}\bigr)\in\widehat{\mathcal{S}}$ be a strict intensity profile and $s$ its canonical representation. Given allocations $x$ and $y$, the former is said to intensity-dominate the latter if for every pair of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$ it holds that $s_{i}(x_{i},x_{j})\geq s_{j}(y_{j},y_{i})$, and the inequality is strict for some pair.
@@
An allocation is intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ if it is Pareto-efficient at the induced preference profile $P_{S}\in\mathcal{P}$ and is not intensity-dominated by another Pareto-efficient allocation at $P_{S}$. It is intensity-dominant if it intensity-dominates every other Pareto-efficient allocation at $P_{S}$.
@@ 4.2 Intensity-Efficient and Utilitarian Allocations Are Distinct
[…]
@@
This potential divergence between intensity efficiency and the hallmark cardinal-welfarist allocation criterion is important because it shows that the hereby proposed notion is indeed distinct even under the richest possible informational assumptions. Furthermore, while perhaps not obvious ex ante, as is clarified by Example 1 and the discussion in Section 3, the reason why such a divergence may occur is that the classical utilitarian criterion is guided by the utility differences across agents. By contrast, in this special environment where intensity relations are defined by the relative ordering rather than the absolute magnitude of utility differences, the Interpersonal Comparisons of Ordinal Intensities postulate upon which the intensity efficiency criterion is built compares the relative ranking of those differences across agents’ intensity orderings.
@@
Example 1 (continued).
Observe that the slightly modified strict utility profile in Example 1 where the second agent’s utility vector is marginally altered from $(46,33,21)$ to $(46,34,20)$ remains strict but now also becomes balanced. Under the modified profile the two Pareto-efficient allocations are both utilitarian and intensity-efficient. $\blacklozenge$